# Momentum accelerated power iterations and the restarted Lanczos method

### v1

@@ 1 Introduction
The Lanczos algorithm and its variants are among the most efficient methods for approximating an extremal eigenvector, or extremal set of eigenvectors of a large sparse symmetric or complex Hermitian matrix [aishima15, saad80, saad11-book, urschel21]. Its implementation is generally limited; however, by computational and memory resources, and in practical computing it may be run as a restarted method, as in [sorensen1992implicit, WuSi00]. As shown in [aishima15], global convergence properties are maintained by restarted variants of the method, so long as the initial vector is not orthogonal to the target eigenvector. However, restarting limits the overall efficiency. On the other hand, recently introduced momentum accelerated power iterations, developed in [XHDMR18], based on [polyak1964], further investigated in [RJRH22], and with a dynamic implementation introduced in [APZ24], feature lower memory and ease of implementation for the same problem class. These methods are closely related to Chebyshev iteration techniques as in [GoVa61, GuRo02, knyazev87], and have the advantages of requiring neither orthogonalization (or re-orthogonaliztion) nor the solution of auxiliary eigenproblems. As shown in [GuRo02], Chebyshev-type methods are also less sensitive to roundoff errors.
Here we consider the question of determining regimes in which either restarted Lanczos or momentum accelerated power iterations is preferable in terms of convergence rate for the same number of matrix-vector multiplications. For symmetric problems with eigenvalues $|\lambda_{1}|>|\lambda_{2}|$, and $|\lambda_{k}|\geq|\lambda_{k+1}|$, for $k\geq 2$, we find that the momentum accelerated power iterations achieve faster convergence rates for the same number of matrix-vector multiplies as restarted Lanczos implemented with increasingly large Krylov subspaces as the ratio $|\lambda_{2}/\lambda_{1}|$ tends to $1$. We also find that combining the two techniques by using the momentum methods as a preconditioner for restarted Lanczos can lead to faster convergence than either method alone in situations where restarted Lanczos with larger Krylov subspaces yields faster convergence than momentum accelerated power iterations. This technique is distinct from the extrapolation method applied to Arnoldi in [PoSc21-arnoldi], and may be generalizable to nonsymmetric problems via the more generalized momentum-type methods in [CMP25-d, CMP25-r]. In what follows, we quantify the regimes where each of the methods considered herein yields faster convergence, and discuss the mechanism behind the successful combination of the two techniques as a momentum-preconditioned restarted Lanczos method.
@@ 1.1 Notation
Suppose $A$ is an $n\times n$ real symmetric or complex Hermitian matrix with eigenvalues $\lambda_{1}\ldots\lambda_{n}$ ordered by decreasing magnitude with $\lambda_{1}>|\lambda_{2}|\geq\cdots\geq|\lambda_{n}|$, with corresponding eigenvectors $\phi_{1},\ldots,\phi_{n}$. Note that we assume there is a unique eigenvalue $\lambda_{1}$ of largest magnitude; however, the assumption that $\lambda_{1}>0$ is for notational convenience and clarity of presentation. If the largest magnitude eigenvalue of $A$ were negative, then the results as stated apply to A-A.
@@ 1.2 Algorithms
We next state the static momentum method with momentum parameter $\beta$, as introduced in [XHDMR18].
@@
As shown in [APZ24, XHDMR18], an efficient implementation of algorithm 1.2 requires a good approximation of $\lambda_{2}$ to approximate the optimal choice of $\beta$, which is $\lambda_{2}^{2}/4$. When we use algorithm 1.2 as a preconditioner for restarted Lanczos, a good approximation to this parameter may be available. Otherwise, the dynamic implementation introduced in [APZ24] can be used to approximate $\lambda_{2}$, and hence the optimal momentum parameter $\beta$, by monitoring the residual convergence rate. The dynamic momentum algorithm stated next was shown in [APZ24] to achieve the efficiency of the optimal static method without explicit a priori knowledge of the spectrum.
@@ 2.1 Chebyshev polynomials
The following results on the Chebyshev polynomials will be used throughout the remainder of the discussion. These results can be found, for instance, in [Gautschi, Chapter 2], [SaVi14, Chapter 2], and [saad11-book, Chapter 4], which refers to [cheney66-approx].
@@ 2.2 Momentum accelerated power iteration
The standard power iteration $u_{k+1}=Ax_{k}$, with normalization $x_{k}=u_{k}/\left\|{u_{k}}\right\|$, for some norm $\left\|{\cdot}\right\|$, is easily seen to satisfy $x_{N}=\widetilde{C}_{N}A^{N}x_{0}=\widetilde{C}_{N}\widetilde{p}_{N}(A)x_{0}$, for normalization constant $\widetilde{C}_{N}$, and $\widetilde{p}_{j}(A)=A^{j}$. See, for instance, [QSS07, chapter 5]. A natural way to change the convergence properties of the iteration is to exchange the monomial $\widetilde{p}_{j}(\cdot)$, for some other polynomial $p_{j}(\cdot)$.
@@
As shown in [XHDMR18], the polynomials $p_{N}(x)$ given by the recurrence relation
@@
which by (2.6) optimizes the convergence to the dominant eigenmode for (2.10) with $\beta=\lambda_{2}^{2}/4$. Sharp results for the convergence of each subdominant eigenmode for $0\leq\lambda^{\ast}\leq\lambda_{1}$ can be found in [APZ24], and are summarized below in subsection 2.2.2. These results will be useful in subsection 4, where we consider the use of this iteration as a preconditioner.
@@ 2.2.1 Normalized iteration
which meets our requirement as a power-like iteration, and thanks to (2.13), has an optimized rate of convergence, so long as $\beta$ can be defined appropriately. The static momentum algorithm 1.2 states (2.2.1) in algorithmic form. Next, we review results on the convergence of each eigenmode under (2.2.1), then we review results on a dynamic implementation from [APZ24] that allows the efficient approximation of the optimal constant $\beta=\lambda_{2}^{2}/4$.
@@ 2.2.2 Convergence of each eigenmode
While (2.12) and (2.13) quantify the convergence to the first eigenmode for iteration (2.2.1), another approach to the analysis, in which we consider this iteration as a power iteration applied to an augmented matrix, will allow us to determine the convergence of each subdominant eigenmode to zero. These results will be useful in our discussion on preconditoning. Moreover, this analysis shows how the method converges for any parameter $\beta\in(0,\lambda_{1}^{2}/4)$, and it shows that the method will not converge if $\beta\geq\lambda_{1}^{2}/4$. The following arguments were introduced in [XHDMR18] and further detailed in [APZ24].
@@
is seen to agree with the iteration given by (2.2.1), for $h_{k}=\left\|u_{k}\right\|$. Thus the convergence of each mode follows from the standard analysis of the power iteration. It is sufficient to determine the eigenvalues $\mu_{\lambda}$ of $A_{\beta}$ corresponding to each eigenvalue of $\lambda$ of $A$, which as shown in [APZ24, XHDMR18], are given by
@@
In the case that $\beta=\lambda_{j}^{2}/4$ for some eigenvalue $\lambda_{j}$ of $A$, the augmented matrix $A_{\beta}$ is actually defective. This is described in some detail in [APZ24]. In particular, the optimal (asymptotic) convergence rate is recovered when $\beta=\lambda_{2}^{2}/4$ yielding
@@ 2.2.3 Dynamic implementation
The main issue preventing an efficient implementation of iteration (2.2.1), i.e., algorithm 1.2, with optimal parameter $\beta=\lambda_{2}^{2}/4$, is that $\lambda_{2}$ is generally a priori unknown. An approach for approximating $\lambda_{2}$ using a deflation technique is proposed in [RJRH22]; however, this requires three matrix-vector multiplies per iteration in the deflation stage of the algorithm, and the guarantee of convergence following this algorithm depends on knowledge of $|\lambda_{1}-\lambda_{2}|$ and $|\lambda_{2}-\lambda_{3}|$. A more recent approach for approximating the optimal parameter $\beta$ is given in algorithm 1.3. Here, the static $\beta$ is replaced by a dynamic approximation $\beta_{k}$ at each iteration. The basis of the approximation is inverting the optimal accelerated rate $\rho(r)=r/(1+\sqrt{1-r^{2}})$ as in (2.27), for $r(\rho)$, where $r=|\lambda_{2}/\lambda_{1}|$, yielding
@@
Within the algorithm, $\rho_{k}$ is computed by the ratio of consecutive normed residuals, denoted $\|d_{k}\|$ in algorithm 1.3, inverted via (2.28) to determine $r_{k+1}$, and multiplied by the Rayleigh quotient to approximate $\lambda_{2}$, hence the optimal $\beta$. The essential mechanism by which $\beta_{k}$ starting from initial $\beta_{0}$ (computed from the ratio of residuals from two consecutive power iterations), tends towards the optimal $\beta=\lambda_{2}^{2}/4$, is that $r(\rho)$ is a contraction map. In [APZ24] we show this by a perturbation argument. This argument demonstrates that, when $r$ is close to $1$, the dynamic approximation becomes increasingly stable. Note that $r$ close to $1$ corresponds to the case where the spectral gap is small, which is the case of practical interest, where the power iteration is slow.
@@
[APZ24, Lemma 3] Let $\rho\in(0,1)$ and consider $\varepsilon$ small enough so that $(2\rho\varepsilon+\varepsilon^{2})/(1+\rho^{2})<1$. Let $\rho_{k}=\rho+\varepsilon$ and define $r_{k+1}=2\rho_{k}/(1+\rho_{k}^{2})$, as in iteration (2.2.1) and algorithm 1.3. Then
@@
In the papers introducing dynamic iterations for the more general polynomial acceleration strategies of [CMP25-d, CMP25-r], the map $r(\rho)$ given by (2.28) is shown to be a contraction. Next we show the analogous result for iteration 2.2.1.
@@
A full convergence result for the dynamic momentum algorithm 1.3 and comparison to the static momentum method algorithm 1.2 with optimal $\beta$ is given in [APZ24]. Here, we will follow the conclusions of that paper, and run our numerical comparisons using the dynamic algorithm 1.3 when we are looking at the momentum accelerated power iteration as a standalone method. When used in conjunction with restarted Lanzcos, we found it is more stable to use the output approximation to $\lambda_{2}$ from restarted Lanzcos in the static momentum algorithm 1.2, instead.
@@
For analysis of Lanczos convergence, we rely on the bounds given by [saad80, saad11-book], which produce estimates of how the error decreases for a given spectrum as the Krylov subspace size $m$ increases. Other important analyses of the Lanzcos process include [KuWo89, urschel21], which focus on spectrum-independent uniform convergence estimates based on matrix dimension, and may yield sharper results for general symmetric matrices, but not when the spectral gap is sufficiently small. The spectrum-dependent convergence theory for the Lanczos process is summarized in [saad11-book, Theorem 6.3], see also [saad80, Theorem 1].
@@
[saad11-book, Theorem 6.3]. The angle between the eigenvector $\phi_{i^{\prime}}$ associated with eigenvalue $\lambda_{i^{\prime}}$ and the Krylov space $\mathcal{K}_{m}=\operatorname{span}\{v_{0},Av_{0},\ldots,A^{m-1}v_{0}\}$, satisfies the inequality
@@
$\tan(\phi_{1},\mathcal{K}_{m})\leq\frac{1}{T_{m-1}\left(1+2\varepsilon_{L}\right)}\tan(x_{0},\phi_{1}),~\text{ with }~\varepsilon_{L}\coloneqq\kappa_{1}=\left(\frac{\lambda_{1}-\lambda_{2^{\prime}}}{\lambda_{2^{\prime}}-\lambda_{n^{\prime}}}\right).$ (2.33)
@@
$\left|T_{m-1}\left(1+2\left(\frac{\lambda_{1}-\lambda_{2^{\prime}}}{\lambda_{2^{\prime}}-\lambda_{n^{\prime}}}\right)\right)\right|=\left|T_{m-1}\left(1+2\varepsilon_{L}\right)\right|\geq\frac{1}{2}\left(1+2\sqrt{\varepsilon_{L}}\right)^{-(m-1)}.$
@@
$\left|T_{m-1}\left(1+2\left(\frac{\lambda_{1}-\lambda_{2^{\prime}}}{\lambda_{2^{\prime}}-\lambda_{n^{\prime}}}\right)\right)\right|^{-1}\leq 2\left(1+2\sqrt{\varepsilon_{L}}\right)^{(m-1)}.$ (2.34)
@@ 3 Comparing rates
To determine regimes in which the momentum method converges faster than restarted Lanczos, we would like to determine the Krylov space dimension $m$ such that
@@
Simplifying (3.1) yields
@@
In [APZ24, XHDMR18], and as summarized in subsection 2.2.2, an explicit asymptotic convergence rate for the (static) momentum method with optimal parameter is found, namely

### v2

# arXiv:2511.05364v2[math.NA] 01 Mar 2026

@@ 1 Introduction
The Lanczos algorithm and its variants are among the most efficient methods for approximating an extremal eigenvector, or extremal set of eigenvectors of a large sparse symmetric or complex Hermitian matrix [1, 17, 18, 21]. Its implementation is generally limited, however, by computational and memory resources, and in practical computing it may be run as a restarted method, as in [20, 22]. As shown in [1], global convergence properties are maintained by restarted variants of the method, so long as the initial vector is not orthogonal to the target eigenvector. However, restarting limits the overall efficiency. On the other hand, recently introduced momentum accelerated power iterations, developed in [23], based on [14], further investigated in [16], and with a dynamic implementation introduced in [2], feature lower memory and ease of implementation for the same problem class. These methods differ in that the method of [23] requires knowledge of the second eigenvalue to set the optimal momentum parameter, while the method of [16] approximates this eigenvalue with a preliminary deflation stage, and the method of [2] uses a posteriori information based on computed residuals and a Rayleigh quotient to approximate the optimal parameter. These methods are closely related to Chebyshev iteration techniques as in [9, 10, 11], and have the advantages of requiring neither orthogonalization (or re-orthogonaliztion) nor the solution of auxiliary eigenproblems. As shown in [10], Chebyshev-type methods are also less sensitive to roundoff errors.
Here we consider the question of determining regimes in which either restarted Lanczos or momentum accelerated power iterations are preferable in terms of convergence rate for the same number of matrix-vector multiplications. For symmetric problems with eigenvalues $|\lambda_{1}|>|\lambda_{2}|$, and $|\lambda_{k}|\geq|\lambda_{k+1}|$, for $k\geq 2$, we find that the momentum accelerated power iterations achieve faster convergence rates for the same number of matrix-vector multiplies as restarted Lanczos implemented with increasingly large Krylov subspaces as the ratio $|\lambda_{2}/\lambda_{1}|$ tends to $1$. We also find that combining the two techniques by using the momentum methods as a preconditioner for restarted Lanczos can lead to faster convergence than either method alone in situations where restarted Lanczos with larger Krylov subspaces yields faster convergence than momentum accelerated power iterations. This technique is distinct from the extrapolation method applied to Arnoldi in [13], and may be generalizable to nonsymmetric problems via the more generalized momentum-type methods in [4, 5]. In what follows, we quantify the regimes where each of the methods considered herein yields faster convergence, and discuss the mechanism behind the successful combination of the two techniques as a momentum-preconditioned restarted Lanczos method.
@@ 1.1 Notation
Suppose $A$ is an $n\times n$ real symmetric or complex Hermitian matrix with eigenvalues $\lambda_{1}\ldots\lambda_{n}$, ordered by decreasing magnitude with $\lambda_{1}>|\lambda_{2}|\geq\cdots\geq|\lambda_{n}|$, with corresponding eigenvectors $\phi_{1},\ldots,\phi_{n}$. Note that we assume there is a unique eigenvalue $\lambda_{1}$ of largest magnitude; however, the assumption that $\lambda_{1}>0$ is for notational convenience and clarity of presentation. If the largest magnitude eigenvalue of $A$ were negative, then the results as stated apply to A-A.
@@ 1.2 Algorithms
We next state the static momentum method with momentum parameter $\beta$, as introduced in [23].
@@
As shown in [2, 23], an efficient implementation of algorithm 1.2 requires a good approximation of $\lambda_{2}$ to approximate the optimal choice of $\beta$, which is $\lambda_{2}^{2}/4$. When we use algorithm 1.2 as a preconditioner for restarted Lanczos, a good approximation to this parameter may be available. Otherwise, the dynamic implementation introduced in [2] can be used to approximate $\lambda_{2}$, and hence the optimal momentum parameter $\beta$, by monitoring the residual convergence rate. The dynamic momentum algorithm stated next was shown in [2] to achieve the efficiency of the optimal static method without explicit a priori knowledge of the spectrum.
@@ 2.1 Chebyshev polynomials
The following results on the Chebyshev polynomials will be used throughout the remainder of the discussion. These results can be found, for instance, in [8, Chapter 2], [19, Chapter 2], and [18, Chapter 4], which refers to [3].
@@ 2.2 Momentum accelerated power iteration
The standard power iteration $u_{k+1}=Ax_{k}$, with normalization $x_{k}=u_{k}/\left\|{u_{k}}\right\|$, for some norm $\left\|{\cdot}\right\|$, is easily seen to satisfy $x_{N}=\widetilde{C}_{N}A^{N}x_{0}=\widetilde{C}_{N}\widetilde{p}_{N}(A)x_{0}$, for normalization constant $\widetilde{C}_{N}$, and $\widetilde{p}_{j}(A)=A^{j}$. See, for instance, [15, chapter 5]. A natural way to change the convergence properties of the iteration is to exchange the monomial $\widetilde{p}_{j}(\cdot)$, for some other polynomial $p_{j}(\cdot)$.
@@
As shown in [23], the polynomials $p_{N}(x)$ given by the recurrence relation
@@
which by (2.6) optimizes the convergence to the dominant eigenmode for (2.10) with $\beta=\lambda_{2}^{2}/4$. Sharp results for the convergence of each subdominant eigenmode for $0\leq\lambda^{\ast}\leq\lambda_{1}$ can be found in [2], and are summarized below in subsection 2.2.2. These results will be useful in subsection 4, where we consider the use of this iteration as a preconditioner.
@@ 2.2.1 Normalized iteration
which meets our requirement as a power-like iteration, and thanks to (2.13), has an optimized rate of convergence, so long as $\beta$ can be defined appropriately. The static momentum algorithm 1.2 states (2.2.1) in algorithmic form. Next, we review results on the convergence of each eigenmode under (2.2.1), then we review results on a dynamic implementation from [2] that allows the efficient approximation of the optimal constant $\beta=\lambda_{2}^{2}/4$.
@@ 2.2.2 Convergence of each eigenmode
While (2.12) and (2.13) quantify the convergence to the first eigenmode for iteration (2.2.1), another approach to the analysis, in which we consider this iteration as a power iteration applied to an augmented matrix, will allow us to determine the convergence of each subdominant eigenmode to zero. These results will be useful in our discussion on preconditoning. Moreover, this analysis shows how the method converges for any parameter $\beta\in(0,\lambda_{1}^{2}/4)$, and it shows that the method will not converge if $\beta\geq\lambda_{1}^{2}/4$. The following arguments were introduced in [23] and further detailed in [2].
@@
is seen to agree with the iteration given by (2.2.1), for $h_{k}=\left\|u_{k}\right\|$. Thus the convergence of each mode follows from the standard analysis of the power iteration. It is sufficient to determine the eigenvalues $\mu_{\lambda}$ of $A_{\beta}$ corresponding to each eigenvalue of $\lambda$ of $A$, which as shown in [2, 23], are given by
@@
In the case that $\beta=\lambda_{j}^{2}/4$ for some eigenvalue $\lambda_{j}$ of $A$, the augmented matrix $A_{\beta}$ is actually defective. This is described in some detail in [2]. In particular, the optimal (asymptotic) convergence rate is recovered when $\beta=\lambda_{2}^{2}/4$ yielding
@@ 2.2.3 Dynamic implementation
The main issue preventing an efficient implementation of iteration (2.2.1), i.e., algorithm 1.2, with optimal parameter $\beta=\lambda_{2}^{2}/4$, is that $\lambda_{2}$ is generally a priori unknown. An approach for approximating $\lambda_{2}$ using a deflation technique is proposed in [16]; however, this requires three matrix-vector multiplies per iteration in the deflation stage of the algorithm, and the guarantee of convergence following this algorithm depends on knowledge of $|\lambda_{1}-\lambda_{2}|$ and $|\lambda_{2}-\lambda_{3}|$. A more recent approach for approximating the optimal parameter $\beta$ is given in algorithm 1.3. Here, the static $\beta$ is replaced by a dynamic approximation $\beta_{k}$ at each iteration. The basis of the approximation is inverting the optimal accelerated rate $\rho(r)=r/(1+\sqrt{1-r^{2}})$ as in (2.27), for $r(\rho)$, where $r=|\lambda_{2}/\lambda_{1}|$, yielding
@@
Within the algorithm, $\rho_{k}$ is computed by the ratio of consecutive normed residuals, denoted $\|d_{k}\|$ in algorithm 1.3, inverted via (2.28) to determine $r_{k+1}$, and multiplied by the Rayleigh quotient to approximate $\lambda_{2}$, hence the optimal $\beta$. The essential mechanism by which $\beta_{k}$ starting from initial $\beta_{0}$ (computed from the ratio of residuals from two consecutive power iterations), tends towards the optimal $\beta=\lambda_{2}^{2}/4$, is that $r(\rho)$ is a contraction map. In [2] we show this by a perturbation argument. This argument demonstrates that, when $r$ is close to $1$, the dynamic approximation becomes increasingly stable. Note that $r$ close to $1$ corresponds to the case where the spectral gap is small, which is the case of practical interest, where the power iteration is slow.
@@
[2, Lemma 3] Let $\rho\in(0,1)$ and consider $\varepsilon$ small enough so that $(2\rho\varepsilon+\varepsilon^{2})/(1+\rho^{2})<1$. Let $\rho_{k}=\rho+\varepsilon$ and define $r_{k+1}=2\rho_{k}/(1+\rho_{k}^{2})$, as in iteration (2.2.1) and algorithm 1.3. Then
@@
In the papers introducing dynamic iterations for the more general polynomial acceleration strategies of [4, 5], the map $r(\rho)$ given in this case by (2.28) is shown to be a contraction. Next we show the analogous result for iteration 2.2.1.
@@
A full convergence result for the dynamic momentum algorithm 1.3 and comparison to the static momentum method algorithm 1.2 with optimal $\beta$ is given in [2]. Here, we will follow the conclusions of that paper, and run our numerical comparisons using the dynamic algorithm 1.3 when we are looking at the momentum accelerated power iteration as a standalone method. When used in conjunction with restarted Lanzcos, we found it is more stable to use the output approximation to $\lambda_{2}$ from restarted Lanzcos in the static momentum algorithm 1.2, instead.
@@
For analysis of Lanczos convergence, we rely on the bounds given by [17, 18], which produce estimates of how the error decreases for a given spectrum as the Krylov subspace size $m$ increases. Other important analyses of the Lanzcos process include [12, 21], which focus on spectrum-independent uniform convergence estimates based on matrix dimension, and may yield sharper results for general symmetric matrices, but not when the spectral gap is sufficiently small. The spectrum-dependent convergence theory for the Lanczos process is summarized in [18, Theorem 6.3], see also [17, Theorem 1].
@@
[18, Theorem 6.3]. The angle between the eigenvector $\phi_{i^{\prime}}$ associated with eigenvalue $\lambda_{i^{\prime}}$ and the Krylov space $\mathcal{K}_{m}=\operatorname{span}\{v_{0},Av_{0},\ldots,A^{m-1}v_{0}\}$, satisfies the inequality
@@
$\tan(\phi_{1},\mathcal{K}_{m})\leq\frac{1}{T_{m-1}\left(1+2\varepsilon_{L}\right)}\tan(x_{0},\phi_{1}),~\text{ with }~\varepsilon_{L}\coloneqq\left(\frac{\lambda_{1}-\lambda_{2^{\prime}}}{\lambda_{2^{\prime}}-\lambda_{n^{\prime}}}\right).$ (2.33)
@@
$\left|T_{m-1}\left(1+2\left(\frac{\lambda_{1}-\lambda_{2^{\prime}}}{\lambda_{2^{\prime}}-\lambda_{n^{\prime}}}\right)\right)\right|=\left|T_{m-1}\left(1+2\varepsilon_{L}\right)\right|\geq\frac{1}{2}\left(1+2\sqrt{\varepsilon_{L}}\right)^{(m-1)}.$
@@
$\left|T_{m-1}\left(1+2\left(\frac{\lambda_{1}-\lambda_{2^{\prime}}}{\lambda_{2^{\prime}}-\lambda_{n^{\prime}}}\right)\right)\right|^{-1}\leq 2\left(1+2\sqrt{\varepsilon_{L}}\right)^{-(m-1)}.$ (2.34)
@@ 3 Comparing rates
To determine regimes in which the momentum method with optimal parameter $\beta=\lambda_{2}^{2}/4$ converges faster than restarted Lanczos, we would like to determine the Krylov space dimension $m$ such that the momentum accelerated convergence rate from (2.13) matches the Lanczos($m$) convergence rate determined by (2.34), that is
@@
Since the dynamic momentum algorithm 1.3, which does not require knowledge of $\lambda_{2}$, has been shown to converge at the same rate, the results here apply to both methods. Simplifying (3.1) yields
@@
In [2, 23], and as summarized in subsection 2.2.2, an explicit asymptotic convergence rate for the (static) momentum method with optimal parameter is found, namely